OpenStudy (anonymous):

A common laboratory preparation for oxygen is 2KCLO3(s)--> 2KCl(s) +3O2 if you were designing an experiment to generate four bottles e(each containing 250.mL) of O2 at 25°C and 762 torr and allowing for 25% waste, what mass of potassium chlorate would be required? Please help I keep failing to solve it

OpenStudy (aaronq):

all chemical reactions must be dealt with in terms of moles. so find the moles of oxygen present in the bottles (you can disregard the volume of 1 of the bottle because you're allowing for a 25% loss). you can do that by using the ideal gas equation: PV=nRT After, you need to find the moles of KCl needed for the reaction. Use the stoichiometric coefficients to find moles produced. Set up a ratio using the species of interest, like so: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients , \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) From here you can isolate what you need. For example: if you have 2 moles of B, how many moles of C can you produce? solve algebraically: \(\dfrac{2}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\rightarrow n_C=\dfrac{2*\color{green}{c}}{\color{blue}{b}}\) -------------------------------------------------------- To convert moles to mass, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.

OpenStudy (anonymous):

Thank you so much! I have a test tomorrow and you really helped!

OpenStudy (aaronq):

no problem, dude !